Use the definition of 0 to show that 5n^5 +4n^4 +
Answer #1

2.5n5 < 5n5 +4n4 +3n3 + 2n2 + n < 5n5 +4n5 +3n5 + 2n5 +  n5

i.e. 5n5 < 5n5 +4n4 +3n3 + 2n2 +  n < 15n5

i.e.5n5 +4n4 +3n3 + 2n2 +  n belongs to Theta of n3

3. n2 <2n2 - n+  3 < 2n2 - n2 +3n2

i.e.   n2 <2n2 - n+  3 < 4n2

i.e  2n2 - n+  3 belongs to Theta of n2

4.

  

Using Big O definition:

f = O(g) iff exist c, n0 > 0 such that forall n >= n0 then 0 <= f(n) <= cg(n)

g = O(h) iff exist k, n1 > 0 such that forall n >= n1 then 0 <= g(n) <= kh(n)

Now take the last unequality and divide all members by c: 0 <= f(n)/c <= g(n) and we can substitute g(n) in the second inequality: 0 <= f(n)/c <= kh(n). Finally multiply all members byc and you obtain 0 <= f(n) <= kch(n) that is the definition of f = O(h):

f = O(h) iff exist j, n2 > 0 such that forall n >= n2 then 0 <= f(n) <= jh(n)

In our case it is: n2 = max(n0, n1) and j = ck.